OpenStudy (anonymous):

A motor scooter purchased for $1,000 depreciates at an annual rate of 15%. Write an exponential function, and graph the function. Use the graph to predict when the value will fall below $100.

OpenStudy (anonymous):

A motor scooter purchased for $1,000 depreciates at an annual rate of 15%. Write an exponential function, and graph the function. Use the graph to predict when the value will fall below $100.

OpenStudy (zpupster):

f(t) = a(1 – r)t Substitute 1,000 for a and 0.15 for r. f(t) = 1000(1 – 0.15)t f(t) = 1000(0.85)t